determine the length of x=4(3+y)^2, 1<=y<=4


1
Expert's answer
2021-01-13T15:55:57-0500

x = 4(3 + y)2

"\implies" dx /dy = 4 . 2(3 + y) = 24 + 8y

1+ (dx /dy)2 = 1 + (24 + 8y)2

arc length = "\int"14 "\sqrt{1+ (dx /dy)^2}" dy = "\int"14 "\sqrt{1+ (24+8y)^2}" dy

substituting u = 24 + 8y

"\implies" du = 8dy , when y =1, u =32 and when y = 4 , u = 56


"\implies" arc length = "\int"456 "\sqrt{1+ (u)^2}" du /8

= [(u"\sqrt{1+ (u)^2}" + ln(u +"\sqrt{1+ (u)^2}")/16​​​] 456

=[(56"\sqrt{1+ (56)^2}" + ln(56 +"\sqrt{1+ (56)^2}")/16​​​]-[(32"\sqrt{1+ (32)^2}" + ln(32 +"\sqrt{1+ (32)^2}")/16​​​]

=[(56"\sqrt{3137}" + ln(56 +"\sqrt{3137}")/16​​​]-[(32"\sqrt{1025}" + ln(32 +"\sqrt{1025}")/16​​​]

=132.0349